Adamantine Architectonics. Adamantine Architectonics consists of a U.S. parent and wholly owned subsidiaries in Malaysia​ (A-Malaysia) and Mexico​ (A-Mexico). Selected portions of their​ non-consolidated balance​ sheets, translated into U.S.​ dollars, are shown in the popup​ window,

A-Malaysia (accounts in ringgits) A-Mexico (accounts in pesos)

Long-term debt 12,300,000 Long-term debt 30,000,000

Shareholders' equity 16,810,000 Shareholders' equity 93,000,000

Adamantine Architectonics (Nonconsolidated Balance Sheet-Selected /Items Only)

Investment in subsidiaries Parent long-term debt 18,000,000

In A-Malaysia 4,100,000 Common stock 6,000,000

In A-Mexico 6,200,000 Retained earnings 25,000,000

Current exchange rates

Malaysia (RM/$) 4.10   

Mexico (PS/$) 15.00

1. What is the debt proportion in​ Adamantine's consolidated balance​ sheet? % ​(Round to two decimal​ places.)
Answer #1

First of all we will find the total debt in consolidated balance sheet of Adamantine's.

Long term debt of Parent : $18,000,000

Add : Long term debt of A-Malaysia in USD = 12,300,000/4.10=$3,000,000

Add : Long term debt of A-Mexico in USD = 30,000,000/15=$2,000,000

Total Consolidated Long term debt = $23,000,000

Consolidated Equity Amount = $6,000,000 + $25,000,000 = $31,000,000

So, the debt proportion in consolidated balance sheet = 23,000,000/31,000,000 = 0.74